### v1

Abstract
The Golub-Welsch algorithm [ Math. Comp., 23: 221-230 (1969)] for computing Gaussian quadrature rules is of importance in estimating quadratic forms. Quadrature rules based on this algorithm have long been assumed to be symmetric. Recent research indicates that the presence of asymmetric quadrature nodes may be more often. Such a divergence has led to varying error analyses of the Lanczos quadrature method. Since symmetry often implies simplicity, it is of great interest to ask when do Lanczos iterations generate symmetric quadrature rules. This paper derives a sufficient condition that ensures symmetric quadrature nodes which partially answers the question that when the Ritz values of a symmetric matrix are symmetrically distributed. Additionally, we establish both lower and upper bounds on the disparity between the minimum Lanczos iterations required for symmetric and asymmetric quadrature.
keywords:
Quadratic forms, Lanczos quadrature method, symmetry of quadrature nodes, Ritz values
{AMS}
65C05, 65D32, 65F15, 65F60, 65G99, 65Y20, 68Q10, 68Q87
1 Introduction
Given a smooth function $f$ defined on $[a,b]$ and a non-negative weight function $\omega$, the Golub-Welsch algorithm [24] can produce a Gaussian quadrature rule,
[…]
where quadrature nodes $\{\theta_{k}\}_{k=1}^{m}$ are the eigenvalues of the tridiagonal Jacobi matrix produced in the Lanczos iteration [30], and the quadrature weights are the squares of the first element of each eigenvector of the Jacobi matrix. Such a Gauss quadrature rule based on the Lanczos method (named Lanczos quadrature method) plays a central role in estimating quadratic forms involving matrix functions
$Q(\boldsymbol{u},f,A)=\boldsymbol{u}^{T}f(A)\boldsymbol{u}\approx\Arrowvert\boldsymbol{u}\Arrowvert_{2}^{2}\sum_{k=1}^{m}\tau_{k}f(\theta_{k})$ (1)
where $A$ is a symmetric matrix, $\boldsymbol{u}$ is a vector, see [23, 21] or the following sessions for more details. The utilization of this method was previously employed in the analysis of iterative methods [20, 35, 4, 9] [36, p.195] and the finite element method [42]. Bai, Golub, and Fahey applied such a method within the realm of Quantum Chromodynamics (QCD) [1], and increasing more applications are emerging. Notably, it can be used to measure the centrality of complex networks. For instance, subgraph centrality can be quantified as $\boldsymbol{e}_{i}^{T}e^{\beta A}\boldsymbol{e}_{i}$ [18], while resolvent-based subgraph centrality is associated with $\boldsymbol{e}_{i}^{T}(I-\alpha A)^{-1}\boldsymbol{e}_{i}$ [17]. Fast computing such centrality is imperative in practical scenarios such as urban public transport [5, 6]. In genomics, statisticians often require approximations of the distribution of quadratic forms (1) with normal distributed vectors [34, 12, 33]. In Gaussian process regression, hyperparameter optimization is achieved through the maximization of a likelihood function, which necessitates the estimation of quadratic forms (1) [13, 44, 53, 54, 52]. The Lanczos quadrature method offers effective computational means for accurately computing these quantities. Furthermore, this method can be integrated into Krylov-spectral methods for solving time-dependent partial differential equations [28, 29].
For a long time, the quadrature nodes generated by the Lanczos algorithm were assumed to be symmetric. For example, the error analysis given by Ubaru, Chen and Saad relied upon such an assumption [48, Section 4.1]. Until very recently, Cortinovis and Kressner highlighted that the quadrature nodes may more frequently exhibit asymmetry, consequently deriving a corresponding error analysis [14][32]. From the standpoint of computational complexity, symmetry typically represents a desirable property due to lower complexity, as evidenced in works such as [38], with further analysis provided in Section 4. Therefore, an intriguing inquiry arises regarding the circumstances under which the Lanczos algorithm would yield symmetric quadrature nodes. This question is equivalent to the following problem.
Problem 1.1.
Given a symmetric matrix $A$ and an initial vector $\boldsymbol{u}$, when are the Ritz values (eigenvalues of the tridiagonal Jacobi matrix obtained in the Lanczos process) symmetrically distributed?
To our knowledge, numerous studies have investigated the spatial distribution [10, 11, 8, 37], convergence characteristics [47, 16], and stabilization techniques [39, 40, 41] concerning Ritz values within the Lanczos method and the Arnoldi process, but none of them directly address our problem. This paper aims at partially answering 1.1 and proposes a sufficient condition ensuring the presence of symmetric Ritz values within Lanczos iterations.
Besides, this paper also analyzes the lower and upper bounds of the difference between the two least required Lanczos iterations $m$ that guarantee the same approximation accuracy of quadratic forms (1). This analysis is based on [48, Theorem 4.2] and [14, Corollary 3] for the cases of symmetric and asymmetric quadrature nodes respectively.
We commence by revisiting the Lanczos quadrature method’s process for estimating quadratic forms (1) in Section 2. In Section 3, we introduce a sufficient condition Theorem 3.16 tailored for the Lanczos quadrature with symmetric quadrature nodes. Subsequently, in Section 4, Corollary 4.3 scrutinizes the potential savings in iterations through such symmetry. Numerical experiments are given in Section 5 to illustrate the validity of our theoretical results, followed by conclusions in Section 6.
2 The Lanczos quadrature method
For a symmetric (positive definite) matrix $A$, its quadratic form $Q(\boldsymbol{u},f,A)$ (1) with vector $\boldsymbol{u}$ and matrix function $f$ can be estimated by the Lanczos quadrature method [23, 21]. The method begins with the eigen-decomposition $A=Q\Lambda Q^{T}$, which leads to $f(A)=Qf(\Lambda)Q^{T}$. One may assume $\Lambda=\mathrm{diag}(\lambda_{1},\lambda_{2},\cdots,\lambda_{n})$ and $\lambda_{\min}=\lambda_{1}\leq\lambda_{2}\leq\ldots\leq\lambda_{n}=\lambda_{\max}$ without loss of generality. Then $f(\Lambda)$ is a diagonal matrix with entries $\{f(\lambda_{j})\}_{j=1}^{n}$. Let $\boldsymbol{u}$ be normalized as $\boldsymbol{v}=\boldsymbol{u}/\Arrowvert\boldsymbol{u}\Arrowvert_{2}$, and $\boldsymbol{\mu}=Q^{T}\boldsymbol{v}$, (1) further reads
$Q(\boldsymbol{u},f,A)=\boldsymbol{u}^{T}f(A)\boldsymbol{u}=\Arrowvert\boldsymbol{u}\Arrowvert_{2}^{2}\boldsymbol{v}^{T}Qf(\Lambda)Q^{T}\boldsymbol{v}=\Arrowvert\boldsymbol{u}\Arrowvert_{2}^{2}\boldsymbol{\mu}^{T}f(\Lambda)\boldsymbol{\mu}.$ (2)
The last quadratic form in (2) can be reformulated as a Riemann-Stieltjes integral $\mathcal{I}$,
[…]
where $\mu_{j}$ is the $j^{th}$ element of $\boldsymbol{\mu}$ and measure $\mu(t)$ is a piecewise constant function of $\mathcal{I}$
$\mu(t)=\left\{\begin{aligned} &\quad 0,&{\rm if}\ t<\lambda_{1}=a,\\ &\sum_{j=1}^{k-1}\mu_{j}^{2},&{\rm if}\ \lambda_{k-1}\leq t<\lambda_{k},k=2,...,n,\\ &\sum_{j=1}^{n}\mu_{j}^{2},&{\rm if}\ t\geq\lambda_{n}=b.\end{aligned}\right.$ (4)
According to the Gauss quadrature rule [22, Chapter 6.2], the Riemann Stieltjes integral $\mathcal{I}$ can be approximated by an $m$-point Lanczos quadrature rule $\mathcal{I}_{m}$ so the last term in equation (3) reads
$\Arrowvert\boldsymbol{u}\Arrowvert_{2}^{2}\mathcal{I}\approx\Arrowvert\boldsymbol{u}\Arrowvert_{2}^{2}\mathcal{I}_{m}=\Arrowvert\boldsymbol{u}\Arrowvert_{2}^{2}\sum_{k=1}^{m}\tau_{k}f(\theta_{k})\equiv Q_{m}(\boldsymbol{u},f,A),$ (5)
where $m$ is the number of quadrature nodes. The quadrature nodes $\{\theta_{k}\}_{k=1}^{m}$ and weights $\{\tau_{k}\}_{k=1}^{m}$ can be obtained by the Lanczos algorithm [1, 22]. Let $T_{m}$ be the Jacobi matrix obtained in the Lanczos process, then the nodes are the eigenvalues of $T_{m}$, and weights are the squares of the first elements of the corresponding normalized eigenvectors [24]. Algorithm 1 outlines how to compute the quadratic form (1) via the Lanczos quadrature method [22, Section 7.2].
Algorithm 1 Lanczos Quadrature Method for Quadratic Form Estimation
Input: Symmetric (positive definite) matrix $A\in\mathbb{R}^{n\times n}$, vector $\boldsymbol{u}\in\mathbb{R}^{n}$, matrix function $f$, steps of Lanczos iterations $m$.
Output: Approximation of the quadratic form $Q_{m}(\boldsymbol{u},f,A)\approx\boldsymbol{u}^{T}f(A)\boldsymbol{u}$.
1:$\boldsymbol{v}_{1}=\boldsymbol{v}=\boldsymbol{u}/\Arrowvert\boldsymbol{u}\Arrowvert_{2}$
2:$\alpha_{1}={\boldsymbol{v}_{1}}^{T}A\boldsymbol{v}_{1}$
3:$\boldsymbol{u}_{2}=A\boldsymbol{v}_{1}-\alpha_{1}\boldsymbol{v}_{1}$
4:for $k=2\text{ to }m$ do
5:     $\beta_{k-1}=\Arrowvert\boldsymbol{u}_{k}\Arrowvert_{2}$
6:     $\boldsymbol{v}_{k}=\boldsymbol{u}_{k}/\beta_{k-1}$
7:     $\alpha_{k}={\boldsymbol{v}_{k}}^{T}A\boldsymbol{v}_{k}$
8:     $\boldsymbol{u}_{k+1}=A\boldsymbol{v}_{k}-\alpha_{k}\boldsymbol{v}_{k}-\beta_{k-1}\boldsymbol{v}_{k-1}$
9:end for
10:$T_{m}=\left[\begin{array}[]{ccccc}\alpha_{1}&\beta_{1}&0&\cdots&0\\ \beta_{1}&\alpha_{2}&\beta_{2}&\ddots&\vdots\\ 0&\beta_{2}&\ddots&\ddots&0\\ \vdots&\ddots&\ddots&\ddots&\beta_{m-1}\\ 0&\cdots&0&\beta_{m-1}&\alpha_{m}\end{array}\right].$
11:$[V,D]=\texttt{eig}(T_{m})$
12:$[\tau_{1},\ldots,\tau_{m}]=\boldsymbol{e}_{1}^{T}V$
13:$[\theta_{1},\ldots,\theta_{m}]^{T}=\texttt{diag}(D)$
14:return $Q_{m}(\boldsymbol{u},f,A)=\Arrowvert\boldsymbol{u}\Arrowvert_{2}^{2}\sum_{k=1}^{m}\tau_{k}f(\theta_{k})$
3 Symmetric Ritz values in the Lanczos iterations
In this section we address the inquiry raised in 1.1. Recall that given a symmetric matrix $A$ and a starting vector $\boldsymbol{u}$, Algorithm 1, executed for $m$ steps, yields a symmetric tridiagonal matrix denoted as $T_{m}$. We are enlightened by the well-known Cauchy interlacing theorem [27, Theorem 4.3.17]. Particularly, [22, Theorem 3.3] details the case of symmetric tridiagonal matrices.
Lemma 3.1.
[22, Theorem 3.3] Let $\{\theta_{k}^{(m)}\}_{k=1}^{m}$ be the eigenvalues of the symmetric tridiagonal matrix $T_{m}$. The eigenvalues of $T_{m+1}$ strictly interlace the eigenvalues of $T_{m}$,
$\theta_{1}^{(m+1)}<\theta_{1}^{(m)}<\theta_{2}^{(m+1)}<\theta_{2}^{(m)}<\cdots<\theta_{m}^{(m)}<\theta_{m+1}^{(m+1)}.$ (6)
It should be noticed that if the eigenvalues of the larger matrix $T_{m+1}$ have successive repeated eigenvalues, e.g., $\theta_{k}^{(m+1)}=\theta_{k+1}^{(m+1)}$, then the interlaced eigenvalue $\theta_{k}^{(m)}$ is equal to these quantities. This observation underscores the validity of employing equal signs within equation (6). Based on Lemma 3.1, we show that the Ritz values generated by the $m$-step Lanczos method with matrix $A$ and vector $\boldsymbol{u}$ are interlaced by the eigenvalues of $A$.
Lemma 3.2.
Let $T_{m}\in\mathbb{R}^{m\times m}$ be the symmetric tridiagonal matrix generated in the $m$-step Lanczos method with symmetric matrix $A\in\mathbb{R}^{n\times n}$ and initial vector $\boldsymbol{u}$. Then under the assumption that values are non-decreasing by index, the eigenvalues of $A$, $\{\lambda_{k}\}_{k=1}^{n}$, interlace the eigenvalues of $T_{m}$, $\{\theta_{k}\}_{k=1}^{m}$,
$\lambda_{k}\leq\theta_{k}\leq\lambda_{k+n-m},k=1,\ldots,m.$ (7)
Proof 3.3.
The proof is completed by two steps. First, Lemma 3.1 helps reveal the relationship of eigenvalues that belong to two tridiagonal Jacobi matrices $T_{m},T_{m+2}$ with order difference two,
$\theta_{k}^{(m+2)}\leq\theta_{k}^{(m)}\leq\theta_{k+2}^{(m+2)},k=1,\ldots,m.$
The inequalities are obtained by first applying Lemma 3.1 on $T_{m},T_{m+1}$ and $T_{m+2}$, and then combining the results. We can further increase the order of the larger symmetric tridiagonal matrix till $n$ and deduce
$\theta_{k}^{(n)}\leq\theta_{k}^{(m)}\leq\theta_{k+n-m}^{(n)},k=1,\ldots,m.$ (8)
Second, the symmetric tridiagonal matrix $T_{n}$ generated by the $n$-step Lanczos method has the same eigenvalues as $A$’s, i.e., $\{\theta_{k}^{(n)}\}_{k=1}^{n}=\{\lambda_{k}\}_{k=1}^{n}$. This is due to the equality
$T_{n}=V_{n}^{T}AV_{n},$
where $V_{n}$ is the orthonormal matrix that stores the vectors $\{\boldsymbol{v}_{k}\}_{k=1}^{n}$ generated during the iteration (see [22, Section 4.1] for more details). Substituting $\{\theta_{k}^{(n)}\}_{k=1}^{n}$ in (8) by $\{\lambda_{k}\}_{k=1}^{n}$ proves (7). The superscript in $\theta_{k}^{(m)}$ is ignored since we would not study the Lanczos method with different numbers of steps in this paper.
Based on this theory, one may deduce that if matrix $A$ possesses an extremely skewed eigenvalue distribution, the resulting Ritz value distribution is unlikely to exhibit symmetry. For example, the spectrum of kernel matrices related to radial basis function, and other linear operators is often skewed distributed and condensed at the origin[50]. Consequently, we regard the symmetry of the eigenvalue distribution as a crucial property for ensuring symmetric Ritz values during the Lanczos iteration. This raises the question:
Problem 3.4.
What types of symmetric matrices $A$ have symmetric eigenvalues?
In Section 3.1 we discuss a type of matrices characterized by the symmetry of eigenvalues, which also find practical applications in various domains.
3.1 Symmetric matrices with symmetric eigenvalues
There are several works demonstrating that the Jordan-Wielandt matrices
$A=\begin{bmatrix}O&B\\ B^{H}&O\end{bmatrix}\in\mathbb{C}^{(n_{1}+n_{2})\times(n_{1}+n_{2})},$ (9)
have symmetric eigenvalue distributions ([19, p. 113] and [31, Section 3.7] briefly described this property without proof when $n_{1}=n_{2}$ and $n_{1}\neq n_{2}$ respectively. Later, [45, Chapter I, Theorem 4.2], [7, Theorem 1.2.2] and [43, Proposition 4.12] detailed such results), where $O$ denotes zero matrix and $B\in\mathbb{C}^{n_{1}\times n_{2}}$.
Such matrices (9) exist in various applications. For instance, the graph of a finite difference matrix is bipartite, meaning that the vertices can be divided into two sets by the red-black order so that no edges exist in each set [25, p. 211] [43, p. 123]. Such matrices can be transformed into the form of Eq. 9 with non-zero diagonal entries. We would utilize this property later in Section 3.3. In the analysis of complex networks, a directed network of $n$ nodes with asymmetric adjacency matrix $B\in\mathbb{R}^{n\times n}$ can be extended to a bipartite undirected network with symmetric block supra-adjacency matrix of form (9) via bipartization [3, 2, 6]. Besides, given a matrix $B\in\mathbb{R}^{n_{1}\times n_{2}}$ and two vectors $\boldsymbol{x}\in\mathbb{R}^{n_{1}},\boldsymbol{y}\in\mathbb{R}^{n_{2}}$, a bilinear form $\boldsymbol{x}^{T}B\boldsymbol{y}$ can be transformed to a quadratic form $\boldsymbol{z}^{T}A\boldsymbol{z}$ with $\boldsymbol{z}=\begin{bmatrix}\boldsymbol{y}\\ \boldsymbol{x}\end{bmatrix}$ and $A=\frac{1}{2}\begin{bmatrix}O&B\\ B^{T}&O\end{bmatrix}$ since
$\boldsymbol{z}^{T}A\boldsymbol{z}=\frac{1}{2}\begin{bmatrix}\boldsymbol{y}^{T}&\boldsymbol{x}^{T}\end{bmatrix}\begin{bmatrix}O&B\\ B^{T}&O\end{bmatrix}\begin{bmatrix}\boldsymbol{y}\\ \boldsymbol{x}\end{bmatrix}=\frac{1}{2}\begin{bmatrix}\boldsymbol{x}^{T}B&\boldsymbol{y}^{T}B^{T}\end{bmatrix}\begin{bmatrix}\boldsymbol{y}\\ \boldsymbol{x}\end{bmatrix}$ $=\frac{\boldsymbol{x}^{T}B\boldsymbol{y}+\boldsymbol{y}^{T}B^{T}\boldsymbol{x}}{2}$
$=\boldsymbol{x}^{T}B\boldsymbol{y}.$
For readers who are new to Jordan-Wielandt matrices and the symmetry of their eigenvalue distributions, please first refer to Lemma 3.5 that exhibits the eigenpairs of (9) when $n_{1}=n_{2}$ [43, Proposition 4.12].
Lemma 3.5.
[43, Proposition 4.12] Let $B\in\mathbb{C}^{n\times n}$ and
$A=\begin{bmatrix}O&B\\ B^{H}&O\end{bmatrix}\in\mathbb{C}^{2n\times 2n}.$
If $\left(\sigma_{i},\begin{bmatrix}\boldsymbol{u}_{i}\\ \boldsymbol{v}_{i}\end{bmatrix}\right)$ is an eigenpair of $B$, then so is $\left(-\sigma_{i},\begin{bmatrix}\boldsymbol{u}_{i}\\ -\boldsymbol{v}_{i}\end{bmatrix}\right)$.
Proof 3.6.
Let $B=U\Sigma V^{H}$ be its singular value decomposition, where
$\Sigma=\mathrm{diag}(\sigma_{1},\sigma_{2},\cdots,\sigma_{n}),$
$U=\begin{bmatrix}\boldsymbol{u}_{1},\cdots,\boldsymbol{u}_{n}\end{bmatrix}$ and $V=\begin{bmatrix}\boldsymbol{v}_{1},\cdots,\boldsymbol{v}_{n}\end{bmatrix}$. Then $B\boldsymbol{v}_{i}=\sigma_{i}\boldsymbol{u}_{i}$ and $B^{H}\boldsymbol{u}_{i}=\sigma_{i}\boldsymbol{v}_{i}$, $i=1,\ldots,n$, followed by
$A\begin{bmatrix}\boldsymbol{u}_{i}\\ \boldsymbol{v}_{i}\end{bmatrix}=\begin{bmatrix}O&B\\ B^{H}&O\end{bmatrix}\begin{bmatrix}\boldsymbol{u}_{i}\\ \boldsymbol{v}_{i}\end{bmatrix}=\begin{bmatrix}B\boldsymbol{v}_{i}\\ B^{H}\boldsymbol{u}_{i}\end{bmatrix}=\begin{bmatrix}\sigma_{i}\boldsymbol{u}_{i}\\ \sigma_{i}\boldsymbol{v}_{i}\end{bmatrix}=\sigma_{i}\begin{bmatrix}\boldsymbol{u}_{i}\\ \boldsymbol{v}_{i}\end{bmatrix},$
$A\begin{bmatrix}\boldsymbol{u}_{i}\\ -\boldsymbol{v}_{i}\end{bmatrix}=\begin{bmatrix}O&B\\ B^{H}&O\end{bmatrix}\begin{bmatrix}\boldsymbol{u}_{i}\\ -\boldsymbol{v}_{i}\end{bmatrix}=\begin{bmatrix}-B\boldsymbol{v}_{i}\\ B^{H}\boldsymbol{u}_{i}\end{bmatrix}=\begin{bmatrix}-\sigma_{i}\boldsymbol{u}_{i}\\ \sigma_{i}\boldsymbol{v}_{i}\end{bmatrix}=-\sigma_{i}\begin{bmatrix}\boldsymbol{u}_{i}\\ -\boldsymbol{v}_{i}\end{bmatrix},$
One may observe that $\left(\sigma_{i},\begin{bmatrix}\boldsymbol{u}_{i}\\ \boldsymbol{v}_{i}\end{bmatrix}\right)$ and $\left(-\sigma_{i},\begin{bmatrix}\boldsymbol{u}_{i}\\ -\boldsymbol{v}_{i}\end{bmatrix}\right),i=1,\ldots,n$ are the eigenpairs of $A$.
For cases of $n_{1}\neq n_{2}$, one could refer to Lemma 3.7 [7] and Corollary 3.8.
Lemma 3.7.
[7, Theorem 1.2.2] Let the singular value decomposition of $B\in\mathbb{C}^{n_{1}\times n_{2}}$ be $B=U\Sigma V^{H},$ where $\Sigma=\begin{bmatrix}\Sigma_{1}&O\\ O&O\end{bmatrix}$, $\Sigma_{1}=\mathrm{diag}(\sigma_{1},\sigma_{2},\cdots,\sigma_{r})$, $r=\mathrm{rank}(B)$,
$U=\left[U_{1},U_{2}\right],U_{1}\in\mathbb{C}^{n_{1}\times r},U_{2}\in\mathbb{C}^{n_{1}\times(n_{1}-r)},$
$V=\left[V_{1},V_{2}\right],V_{1}\in\mathbb{C}^{n_{2}\times r},V_{2}\in\mathbb{C}^{n_{2}\times(n_{2}-r)}.$
Then
$A=\begin{bmatrix}O&B\\ B^{H}&O\end{bmatrix}=P^{H}\begin{bmatrix}\Sigma_{1}&O&O&O\\ O&-\Sigma_{1}&O&O\\ O&O&O&O\\ O&O&O&O\end{bmatrix}P,$
where $P$ is unitary
$P=\frac{1}{\sqrt{2}}\begin{bmatrix}U_{1}&U_{1}&\sqrt{2}U_{2}&O\\ V_{1}&-V_{1}&O&\sqrt{2}V_{2}\end{bmatrix}^{H}.$
Thus $2r$ eigenpairs of $A$ are (±$\left(\pm\sigma_{i},\begin{bmatrix}\boldsymbol{u}_{i}\\ \pm\boldsymbol{v}_{i}\end{bmatrix}\right),i=1,\ldots,r$, and the left eigenvalues 00 repeated $(n_{1}+n_{2}-2r)$ times, where $\boldsymbol{u}_{i}$ and $\boldsymbol{v}_{i}$ are the columns of $U_{1}$ and $V_{1}$ respectively.
Corollary 3.8.
Let the singular value decomposition of $B\in\mathbb{C}^{n_{1}\times n_{2}}$ be $B=U\Sigma V^{H},$ where $\Sigma=\begin{bmatrix}\Sigma_{1}&O\\ O&O\end{bmatrix}$, $\Sigma_{1}=\mathrm{diag}(\sigma_{1},\sigma_{2},\cdots,\sigma_{r})$, $r=\mathrm{rank}(B)$,
$U=\left[U_{1},U_{2}\right],U_{1}\in\mathbb{C}^{n_{1}\times r},U_{2}\in\mathbb{C}^{n_{1}\times(n_{1}-r)},$
$V=\left[V_{1},V_{2}\right],V_{1}\in\mathbb{C}^{n_{2}\times r},V_{2}\in\mathbb{C}^{n_{2}\times(n_{2}-r)}.$
Then the matrix of the form
$A=\begin{bmatrix}\gamma I_{n_{1}}&B\\ B^{H}&\gamma I_{n_{2}}\end{bmatrix}$
has $2r$ eigenpairs of $A$, $\left(\gamma\pm\sigma_{i},\begin{bmatrix}\boldsymbol{u}_{i}\\ \pm\boldsymbol{v}_{i}\end{bmatrix}\right),i=1,\ldots,r$, and the left eigenvalues $\gamma$ repeated $(n_{1}+n_{2}-2r)$ times, where $\boldsymbol{u}_{i}$ and $\boldsymbol{v}_{i}$ are the columns of $U_{1}$ and $V_{1}$ respectively.
Nonetheless, presuming a symmetric eigenvalue distribution may not suffice to ensure the symmetry of Ritz values. Therefore, we proceed to inquire:
Problem 3.9.
Given a symmetric matrix $A$ with symmetric eigenvalues and an initial vector $\boldsymbol{u}$, when are the Ritz values (eigenvalues of the tridiagonal Jacobi matrix obtained in the Lanczos process) symmetrically distributed?
To answer this question, we refer to the case of Jacobi orthogonal polynomials, wherein the distributions of their roots exhibit symmetry when the associated continuous weight functions are even. We conjecture that the symmetry of Ritz values within Lanczos iterations is intricately linked to the underlying discrete measure function $\mu(t)$ as outlined in (4). Elaborate explanations and discussions regarding this conjecture are provided in Section 3.2.
3.2 Central symmetric discrete measure functions
Note that the zeros of the Jacobi polynomials are symmetric about 00 when the corresponding continuous weight function $\omega(t)=(1-t)^{\alpha}(1+t)^{\beta},t\in[-1,1]$ is even and the measure function $\mu(t)$ is central symmetric.
Lemma 3.10.
Let $P_{n}^{\{\alpha,\beta\}}(t)$ be the $n$-degree Jacobi orthogonal polynomial with the weight function $\omega(t)=(1-t)^{\alpha}(1+t)^{\beta}$ on the interval $\left[-1,1\right]$, where $\alpha,\beta>-1$. If $\alpha=\beta$, the zeros of $P_{n}^{\{\alpha,\beta\}}(t)$ will be symmetric about 0.
Proof 3.11.
By the Rodrigues’ formula [46, Chapter IV], the $n^{th}$-degree Jacobi polynomial is written by
$P_{n}^{\{\alpha,\beta\}}(t)=\frac{(-1)^{n}}{2^{n}n!}(1-t)^{-\alpha}(1+t)^{-\beta}\frac{d^{n}}{dt^{n}}\left[(1-t)^{n+\alpha}(1+t)^{n+\beta}\right].$ (10)
If $\alpha=\beta$, the weight function $\omega(t)$ is even and (10) further reads
$P_{n}^{\{\alpha,\alpha\}}(t)=\frac{(-1)^{n}}{2^{n}n!}(1-t^{2})^{-\alpha}\frac{d^{n}}{dt^{n}}\left[(1-t^{2})^{n+\alpha}\right].$
It is trivial that $P_{n}^{\{\alpha,\alpha\}}(t)=(-1)^{n}P_{n}^{\{\alpha,\alpha\}}(-t).$ When $n$ is even, $P_{n}^{\{\alpha,\alpha\}}$ is an even function so the zeros are symmetric about 00. For odd $n$, $P_{n}^{\{\alpha,\alpha\}}(0)=0$, and the rest zeros remain symmetric about 00 due to odd function $P_{n}^{\{\alpha,\alpha\}}$.
Our initial presumption suggests that the weight function associated with Lanczos iterations should exhibit evenness to ensure the symmetry of Ritz values. However, it’s not feasible to explicitly delineate the weight functions for Lanczos orthogonal polynomials generated by the Lanczos method. The only information in hand is the corresponding discrete step measure function Eq. 4 with jumps at the eigenvalues of matrix.
Fortunately, the requirement of an even weight function $\omega(t)$ aligns with the assumption of a centrally symmetric measure function $\mu(t)$ within its domain, where $d\mu(t)=\omega(t)dt$. Consequently, we consider the central symmetry of $\mu(t)$ as another pivotal property essential for substantiating the symmetry of Ritz values in Lanczos iterations.
One may observe that (4) is defined by the components of $\boldsymbol{\mu}=Q^{T}\boldsymbol{v}$, where $Q$ denotes the orthonormal matrix in the eigen-decomposition of $A=Q\Lambda Q^{T}$, and $\boldsymbol{v}=\boldsymbol{u}/\Arrowvert\boldsymbol{u}\Arrowvert_{2}$ represents the normalized initial vector. The term $\mu_{i}^{2}$ signifies the increment of $\mu(t)$ from $t=\lambda_{i}$ to $t=\lambda_{i+1}$, where $\mu_{i}$ denotes the $i^{th}$ element of $\boldsymbol{\mu}$. Building upon our conjecture, we postulate that the first and last $i^{th}$ step increments $\mu_{i}^{2}$ of $\mu(t)$ are equal for $i=1,\ldots,\lfloor n/2\rfloor$. The subsequent definition aids in elucidating this assumption.
Definition 3.12.
A vector $\boldsymbol{w}\in\mathbb{R}^{n}$ is said to be an absolute palindrome if
$|w_{i}|=|w_{n+1-i}|,i=1,\ldots,\lfloor\frac{n}{2}\rfloor.$
Similar to the question raised in 3.4, we would like know,
Problem 3.13.
Given a symmetric matrix $A=Q\Lambda Q^{T}$ and a normalized initial vector $\boldsymbol{v}$. What conditions will guarantee an absolute palindrome $\boldsymbol{\mu}=Q^{T}\boldsymbol{v}$?
Theorem 3.14 establishes the existence of such conditions.
Theorem 3.14.
Let $Q=[\boldsymbol{q}_{1},\ldots,\boldsymbol{q}_{n}]\in\mathbb{R}^{n\times n}$ be an orthonormal matrix and $\boldsymbol{v}\in\mathbb{R}^{n}$ be a normalized initial vector for the Lanczos iteration. Then $\boldsymbol{v}$ can be written as
$\boldsymbol{v}=\sum_{i=1}^{n}\xi_{i}\boldsymbol{q}_{i}$
with real numbers $\{\xi_{i}\}_{i=1}^{n}$. If $\boldsymbol{\xi}=[\xi_{1},\ldots,\xi_{n}]^{T}$ is an absolute palindrome, then $\boldsymbol{\mu}=Q^{T}\boldsymbol{v}$ is also an absolute palindrome.
Proof 3.15.
It is trivial that the columns of $Q=[\boldsymbol{q}_{1},\cdots,\boldsymbol{q}_{n}]$ form a basis of $\mathbb{R}^{n}$, then we may write $\boldsymbol{v}=\xi_{1}\boldsymbol{q}_{1}+\xi_{2}\boldsymbol{q}_{2}+\cdots+\xi_{n}\boldsymbol{q}_{n}$ without loss of generality and $\boldsymbol{\mu}$ reads
$\boldsymbol{\mu}=Q^{T}\boldsymbol{v}=\begin{bmatrix}\boldsymbol{q}_{1}^{T}\\ \boldsymbol{q}_{2}^{T}\\ \vdots\\ \boldsymbol{q}_{n}\end{bmatrix}\left[\xi_{1}\boldsymbol{q}_{1}+\xi_{2}\boldsymbol{q}_{2}+\cdots+\xi_{n}\boldsymbol{q}_{n}\right]=\begin{bmatrix}\xi_{1}\\ \xi_{2}\\ \vdots\\ \xi_{n}\end{bmatrix}=\boldsymbol{\xi}.$
This indicates that the assumption of $|\mu_{i}|=|\mu_{n+1-i}|$ is equivalent to $|\xi_{i}|=|\xi_{n+1-i}|,\\ i=1,\ldots,\lfloor\frac{n}{2}\rfloor$. There must exist certain vector $\boldsymbol{\xi}$ that guarantees the property. Moreover, suppose $n$ is even, for vector $\boldsymbol{\xi}$, it can be expressed as
$\boldsymbol{\xi}=\xi_{1}\begin{bmatrix}1\\ 0\\ \vdots\\ 0\\ 0\\ \vdots\\ 0\\ \pm 1\end{bmatrix}+\xi_{2}\begin{bmatrix}0\\ 1\\ \vdots\\ 0\\ 0\\ \vdots\\ \pm 1\\ 0\end{bmatrix}+\cdots+\xi_{\frac{n}{2}}\begin{bmatrix}0\\ 0\\ \vdots\\ 1\\ \pm 1\\ \vdots\\ 0\\ 0\end{bmatrix}=\xi_{1}\boldsymbol{s}_{1}^{\pm}+\xi_{2}\boldsymbol{s}_{2}^{\pm}+\cdots\xi_{\frac{n}{2}}\boldsymbol{s}_{\frac{n}{2}}^{\pm}.$
Such independent vectors $\{\boldsymbol{s}_{j}^{\pm}\}_{j=1}^{n/2}$ form a subspace $S=\mathrm{span}\{\boldsymbol{s}_{1}^{\pm},\boldsymbol{s}_{2}^{\pm},\ldots,\boldsymbol{s}_{\frac{n}{2}}^{\pm}\}$, where the dimension is $n/2$. Note that there are $2^{\frac{n}{2}}$ such subspaces. For odd $n$, there would be $2^{\frac{n}{2}+1}$ subspaces due to the additional vector $\boldsymbol{s}_{\frac{n}{2}+1}^{\pm}=(0,\cdots,0,\pm 1,0,\cdots,0)^{T}$.
Having established the presence of a symmetric eigenvalue distribution of matrix $A$ and the existence of an absolute palindrome $\boldsymbol{\mu}$, we proceed to the subsequent subsection where we illustrate that these two conditions suffice to ensure symmetric Ritz values in the Lanczos process.
3.3 Sufficient condition of symmetric Ritz values in the Lanczos iterations
In this subsection, we provide a comprehensive elucidation of the sufficient condition that guarantees symmetric quadrature nodes and symmetrically equal weights derived from the Lanczos iteration, as succinctly summarized in Theorem 3.16.
Theorem 3.16.
Let $A=Q\Lambda Q^{T}\in\mathbb{R}^{n\times n}$ be symmetric, $\Lambda=\mathrm{diag}(\lambda_{1},\lambda_{2},\cdots,\lambda_{n})$ with eigenvalues $\lambda_{1}\leq\lambda_{2}\leq\ldots\leq\lambda_{n}$ symmetric about $\bar{\lambda}$, and $\boldsymbol{v}\in\mathbb{R}^{n}$ be a normalized initial vector for the Lanczos iteration. For the $m$-step Lanczos method with $A$ and $\boldsymbol{v}$ ($m=1,\ldots,n$), if vector $\boldsymbol{\mu}=Q^{T}\boldsymbol{v}$ is an absolute palindrome, the distribution of $m$ Ritz values will be symmetric about $\bar{\lambda}$, and the quadrature weights corresponding to the pairs of symmetric quadrature nodes are equal.
The proof of Theorem 3.16 is completed with the aid of Lemma 3.17, which demonstrates the tridiagonal Jacobi matrix $T_{m}$ generated by the $m$-step Lanczos iteration has constant diagonal entries $\bar{\lambda}$.
Lemma 3.17.
Let $A=Q\Lambda Q^{T}\in\mathbb{R}^{n\times n}$ be symmetric, $\Lambda=\mathrm{diag}(\lambda_{1},\lambda_{2},\cdots,\lambda_{n})$ with eigenvalues $\lambda_{1}\leq\lambda_{2}\leq\ldots\leq\lambda_{n}$ symmetric about $\bar{\lambda}$, and $\boldsymbol{v}\in\mathbb{R}^{n}$ be a normalized initial vector for the Lanczos iteration. For the $m$-step Lanczos method with $A$ and $\boldsymbol{v}$ ($m=1,\ldots,n$), if vector $\boldsymbol{\mu}=Q^{T}\boldsymbol{v}$ is an absolute palindrome, the Jacobi matrix $T_{m}$ generated by $m$-step Lanczos iteration will have constant diagonal entries $\bar{\lambda}$.
Proof 3.18.
The proof is divided into two cases: even $n$ and odd $n$. We first study the even case and then extend the conclusion to the odd one. For the Lanczos iteration with initial vector $\boldsymbol{v}_{1}=\boldsymbol{v}$, the values $\alpha_{k}$ are computed as
$\alpha_{k}=\boldsymbol{v}_{k}^{T}A\boldsymbol{v}_{k}=\boldsymbol{v}_{k}^{T}Q\Lambda Q^{T}\boldsymbol{v}_{k}=\boldsymbol{\mu}_{k}^{T}\Lambda\boldsymbol{\mu}_{k},k\geq 1$
where normalized vectors $\boldsymbol{v}_{k}=\boldsymbol{u}_{k}/\Arrowvert\boldsymbol{u}_{k}\Arrowvert$ and $\boldsymbol{\mu}_{k}=Q^{T}\boldsymbol{v}_{k},k\geq 2$. For vectors $\boldsymbol{u}_{k}$, there exists a three-term recurrence between them [22],
$\boldsymbol{u}_{k}=A\boldsymbol{v}_{k-1}-\alpha_{k-1}\boldsymbol{v}_{k-1}-\beta_{k-2}\boldsymbol{v}_{k-2},k\geq 3$
$\boldsymbol{u}_{2}=A\boldsymbol{v}_{1}-\alpha_{1}\boldsymbol{v}_{1}$. Based on the assumption that $\boldsymbol{\mu}_{1}=\boldsymbol{\mu}$ is an absolute palindrome, it can be expressed of the form
$\boldsymbol{\mu}_{1}=\begin{bmatrix}\mathrm{sgn}(\mu_{1,1})|\mu_{1,1}|\\ \mathrm{sgn}(\mu_{1,2})|\mu_{1,2}|\\ \vdots\\ \mathrm{sgn}(\mu_{1,n-1})|\mu_{1,2}|\\ \mathrm{sgn}(\mu_{1,n})|\mu_{1,1}|\end{bmatrix},$ (11)
where $\mu_{1,j}$ (resp. $\mathrm{sgn}(\mu_{1,j})$) denotes the $j^{th}$ component of $\boldsymbol{\mu}_{1}$ (resp. the sign of it), $j=1,\ldots,n$. We propose that all $\boldsymbol{\mu}_{k}$’s are absolute palindromes, $k\geq 1$. More specifically, for odd $k$, $\boldsymbol{\mu}_{k}$ has the form
$\boldsymbol{\mu}_{k}=\begin{bmatrix}\mathrm{sgn}(\mu_{1,1})c_{k,1}|\mu_{1,1}|\\ \mathrm{sgn}(\mu_{1,2})c_{k,2}|\mu_{1,2}|\\ \vdots\\ \mathrm{sgn}(\mu_{1,n-1})c_{k,2}|\mu_{1,2}|\\ \mathrm{sgn}(\mu_{1,n})c_{k,1}|\mu_{1,1}|\end{bmatrix},$ (12)
where $c_{k,j},j=1,\ldots,n/2,$ are the relevant coefficients with respect to the components of $\boldsymbol{\mu}_{k}$. While for even $k$, $\boldsymbol{\mu}_{k}$ has form
$\boldsymbol{\mu}_{k}=\begin{bmatrix}\mathrm{sgn}(\mu_{1,1})c_{k,1}|\mu_{1,1}|\\ \mathrm{sgn}(\mu_{1,2})c_{k,2}|\mu_{1,2}|\\ \vdots\\ -\mathrm{sgn}(\mu_{1,n-1})c_{k,2}|\mu_{1,2}|\\ -\mathrm{sgn}(\mu_{1,n})c_{k,1}|\mu_{1,1}|\end{bmatrix}.$ (13)
This proposition can be proved by mathematical induction. One may observe that if the normalized vector $\boldsymbol{\mu}_{k}$ is in either form (12) or (13), then $\sum_{j=1}^{n/2}\mu_{k,j}^{2}=1/2$. Furthermore, since $A$ has symmetric eigenvalue distribution about $\bar{\lambda}$,
$\alpha_{k}$ $=\boldsymbol{\mu}_{k}^{T}\Lambda\boldsymbol{\mu}_{k}=\sum_{j=1}^{n}\mu_{k,j}^{2}\lambda_{j}$
$=(\lambda_{1}+\lambda_{n})\mu_{k,1}^{2}+(\lambda_{2}+\lambda_{n-1})\mu_{k,2}^{2}+\cdots+(\lambda_{\frac{n}{2}}+\lambda_{\frac{n}{2}+1})\mu_{k,\frac{n}{2}}^{2}$
$=2\bar{\lambda}\sum_{j=1}^{n/2}\mu_{k,j}^{2}=\bar{\lambda}.$
We begin with $k=1$, where $\boldsymbol{\mu}_{1}$ (11) is clearly in the form of (12), so $\alpha_{1}=\bar{\lambda}$.
For $k=2$, $\tilde{\boldsymbol{\mu}}_{2}$ reads
$\tilde{\boldsymbol{\mu}}_{2}=Q^{T}\boldsymbol{u}_{2}=Q^{T}Q\Lambda Q^{T}\boldsymbol{v}_{1}-\alpha_{1}Q^{T}\boldsymbol{v}_{1}=(\Lambda-\bar{\lambda}I)\boldsymbol{\mu}_{1}.$
Let $\tilde{\lambda}_{j}=\lambda_{j}-\bar{\lambda},j=1,\ldots,n$, then
$\tilde{\boldsymbol{\mu}}_{2}=\begin{bmatrix}\mathrm{sgn}(\mu_{1,1})\tilde{\lambda_{1}}|\mu_{1,1}|\\ \mathrm{sgn}(\mu_{1,2})\tilde{\lambda_{2}}|\mu_{1,2}|\\ \vdots\\ -\mathrm{sgn}(\mu_{1,n-1})\tilde{\lambda_{2}}|\mu_{1,2}|\\ -\mathrm{sgn}(\mu_{1,n})\tilde{\lambda_{1}}|\mu_{1,1}|\end{bmatrix},$
since $\tilde{\lambda}_{j}=-\tilde{\lambda}_{n+1-j},j=1,\ldots,n/2$. It is trivial that its normalized vector $\boldsymbol{\mu}_{2}=\tilde{\boldsymbol{\mu}}_{2}/\Arrowvert\tilde{\boldsymbol{\mu}}_{2}\Arrowvert$ has the form Eq. 13, then $\alpha_{2}=\bar{\lambda}$.
Now suppose for positive integer $i$, $\boldsymbol{\mu}_{2i-1}$ (resp. $\boldsymbol{\mu}_{2i}$) has the form (12) (resp. (13)). From this supposition, $\alpha_{2i-1}=\alpha_{2i}=\bar{\lambda}.$ We want to prove for $i+1$ the statement still holds. Based on the three-term recurrence between $\boldsymbol{\mu}_{k}$’s, we derive
$\tilde{\boldsymbol{\mu}}_{2i+1}=Q^{T}\boldsymbol{u}_{2i+1}$ $=Q^{T}Q\Lambda Q^{T}\boldsymbol{v}_{2i}-\alpha_{2i}Q^{T}\boldsymbol{v}_{2i}-\beta_{2i-1}Q^{T}\boldsymbol{v}_{2i-1}$
$=(\Lambda-\bar{\lambda}I)\boldsymbol{\mu}_{2i}-\beta_{2i-1}\boldsymbol{\mu}_{2i-1}$
$=\begin{bmatrix}\mathrm{sgn}(\mu_{1,1})\left(\tilde{\lambda}_{1}c_{2i,1}-\beta_{2i-1}c_{2i-1,1}\right)|\mu_{1,1}|\\ \mathrm{sgn}(\mu_{1,2})\left(\tilde{\lambda}_{2}c_{2i,2}-\beta_{2i-1}c_{2i-1,2}\right)|\mu_{1,2}|\\ \vdots\\ \mathrm{sgn}(\mu_{1,n-1})\left(-(-\tilde{\lambda}_{2})c_{2i,2}-\beta_{2i-1}c_{2i-1,2}\right)|\mu_{1,2}|\\ \mathrm{sgn}(\mu_{1,n})\left(-(-\tilde{\lambda}_{1})c_{2i,1}-\beta_{2i-1}c_{2i-1,1}\right)|\mu_{1,1}|\end{bmatrix},$
which demonstrates that $\boldsymbol{\mu}_{2i+1}$ can also be written as (12), and thus $\alpha_{2i+1}=\bar{\lambda}$. Further, $\boldsymbol{\mu}_{2i+2}$ reads
$\tilde{\boldsymbol{\mu}}_{2i+2}=Q^{T}\boldsymbol{u}_{2i+2}$ $=Q^{T}Q\Lambda Q^{T}\boldsymbol{v}_{2i+1}-\alpha_{2i+1}Q^{T}\boldsymbol{v}_{2i+1}-\beta_{2i}Q^{T}\boldsymbol{v}_{2i}$
$=(\Lambda-\bar{\lambda}I)\boldsymbol{\mu}_{2i+1}-\beta_{2i}\boldsymbol{\mu}_{2i}$
$=\begin{bmatrix}\mathrm{sgn}(\mu_{1,1})\left(\tilde{\lambda}_{1}c_{2i+1,1}-\beta_{2i}c_{2i,1}\right)|\mu_{1,1}|\\ \mathrm{sgn}(\mu_{1,2})\left(\tilde{\lambda}_{2}c_{2i+1,2}-\beta_{2i}c_{2i,2}\right)|\mu_{1,2}|\\ \vdots\\ \mathrm{sgn}(\mu_{1,n-1})\left(-\tilde{\lambda}_{2}c_{2i+1,2}-(-\beta_{2i}c_{2i,2})\right)|\mu_{1,2}|\\ \mathrm{sgn}(\mu_{1,n})\left(-\tilde{\lambda}_{1}c_{2i+1,1}-(-\beta_{2i}c_{2i,1})\right)|\mu_{1,1}|\end{bmatrix},$
which is of the form (13) so $\alpha_{2i+2}=\bar{\lambda}$. The proof of even case ends.
When $n$ is odd, for any $k=1,\ldots,m$, $\alpha_{k}$ reads
$\alpha_{k}=\boldsymbol{\mu}_{k}^{T}\Lambda\boldsymbol{\mu}_{k}=\sum_{j=1}^{\lfloor\frac{n}{2}\rfloor}\mu_{k,j}^{2}\lambda_{j}+\mu_{k,\lfloor\frac{n}{2}\rfloor+1}^{2}\bar{\lambda}+\sum_{j=\lfloor\frac{n}{2}\rfloor+2}^{n}\mu_{k,j}^{2}\lambda_{j}.$
The only difference lays in a single term (i.e., $j=\lfloor n/2\rfloor+1$) that has no pair like every component in even case does. However, one may follow the proof we presented for even $n$ to deduce that the two summations are still equal here and $\Arrowvert\boldsymbol{\mu}_{k}\Arrowvert=1$, resulting in
$\alpha_{k}$ $=(\lambda_{1}+\lambda_{n})\mu_{k,1}^{2}+(\lambda_{2}+\lambda_{n-1})\mu_{k,2}^{2}+\cdots+(\lambda_{\lfloor\frac{n}{2}\rfloor}+\lambda_{\lfloor\frac{n}{2}\rfloor+2})\mu_{k,\lfloor\frac{n}{2}\rfloor}^{2}+\bar{\lambda}\mu_{k,\lfloor\frac{n}{2}\rfloor+1}^{2}$
$=2\bar{\lambda}(1-\mu_{k,\lfloor\frac{n}{2}\rfloor+1}^{2})/2+\bar{\lambda}\mu_{k,\lfloor\frac{n}{2}\rfloor+1}^{2}=\bar{\lambda},$
so $\alpha_{k}=\bar{\lambda},k\geq 1$ for both odd and even $n$. Then, the Jacobi matrix generated by the $m$-step Lanczos iteration would have constant diagonal entries $\bar{\lambda}$, where $m=1,\ldots,n$.
3.4 Proof of main results
It is worth noting that every symmetric tridiagonal matrix with constant diagonal elements $\bar{\lambda}$ can be rearranged into the form
$\begin{bmatrix}\bar{\lambda}I_{n_{1}}&B\\ B^{H}&\bar{\lambda}I_{n_{2}}\end{bmatrix}.$ (14)
This rearrangement follows the well-known red-black ordering [25, p. 211] [43, p. 123], where either $n_{1}=n_{2}$ or $n_{1}=n_{2}+1$. According to Corollary 3.8, if the conditions specified in Theorem 3.16 are satisfied, the Jacobi matrix $T_{m}$ yielded by the $m$-step Lanczos iteration will possess a symmetric eigenvalue distribution, signifying that the $m$ Ritz values exhibit symmetry about $\bar{\lambda}$. Furthermore, owing to the equivalence of the pairs of first entries in the eigenvectors corresponding to the symmetric eigenvalues (as demonstrated in Lemma 3.5, which also applies to Lemma 3.7 and Corollary 3.8), one can ascertain symmetrically equal quadrature weights for such Gaussian quadrature rules. The proof of Theorem 3.16 is thereby concluded.
In Section 5.1, we proceed to demonstrate the validity of Theorem 3.16.
4 Computation complexity for symmetric and asymmetric quadrature nodes
Given a symmetric (positive definite) matrix $A$, a normalized starting vector $\boldsymbol{v}$ and error tolerance $\epsilon$, the theoretical lower bounds of the Lanczos iteration $m$ that guarantee
$|Q(\boldsymbol{v},f,A)-Q_{m}(\boldsymbol{v},f,A)|=|\mathcal{I}-\mathcal{I}_{m}|\leq\epsilon$ (15)
are different for the case of symmetric quadrature nodes and the asymmetric one (see [48, Section 4.1] and [14, Section 3]). Note that the error of $|\mathcal{I}-\mathcal{I}_{m}|$ by [48, 14] is analyzed based on a Bernstein ellipse $E_{\rho}$ with foci ±$\pm 1$ and elliptical radius $\rho>1$, which is formed via an analytic continuation admitted by function $g:[-1,1]\to\mathbb{R}$, while an affine transform $h:[-1,1]\to[\lambda_{\min},\lambda_{\max}]$ is required to ensure $g=f(h(\cdot))$. In detail, the radius $\rho$ is a function of matrix $A$’s condition number $\kappa=\lambda_{\max}/\lambda_{\min}>1$.
Since the choice of $m$ determines the computational complexity $\mathcal{O}(n^{2}m)$, it is imperative to scrutinize the disparity between the two theoretical bounds of $m$ proposed by Lemma 4.1 [48, Theorem 4.2] and Lemma 4.2 [14, Corollary 3] for symmetric and asymmetric Ritz values. To commence, let us provide a brief overview of these theories.
Lemma 4.1.
[48, Theorem 4.2] Let $g$ be analytic in $\left[-1,1\right]$ and analytically continuable in the open Bernstein ellipse $E_{\rho}$ with foci ±$\pm 1$ and elliptical radius $\rho>1$. Let $M_{\rho}$ be the maximum of $|g(t)|$ on $E_{\rho}$. Then the $m$-point Lanczos quadrature approximation satisfies
$|\mathcal{I}-\mathcal{I}_{m}|\leq\frac{4M_{\rho}}{1-\rho^{-2}}\rho^{-2m}.$ (16)
Lemma 4.2.
[14, Corollary 3] Let $g$ be analytic in $\left[-1,1\right]$ and analytically continuable in the open Bernstein ellipse $E_{\rho}$ with foci ±$\pm 1$ and elliptical radius $\rho>1$. Let $M_{\rho}$ be the maximum of $|g(t)|$ on $E_{\rho}$. Then the $m$-point Lanczos quadrature approximation with asymmetric quadrature nodes satisfies
$|\mathcal{I}-\mathcal{I}_{m}|\leq\frac{4M_{\rho}}{1-\rho^{-1}}\rho^{-2m}.$ (17)
Note that the elliptical radius $\rho$ is not fixed. In this paper we set
$\rho=\frac{\lambda_{\max}+\lambda_{\min}}{\lambda_{\max}-\lambda_{\min}}=\frac{\sqrt{\kappa}+1}{\sqrt{\kappa}-1}.$ (18)
Then combining (15), (17) and (16) we have
$m_{\text{asym}}\geq\frac{1}{2\log(\rho)}\cdot\left[\log(4M_{\rho})-\log(1-\rho^{-1})-\log(\epsilon)\right],$ (19)
$m_{\text{sym}}\geq\frac{1}{2\log(\rho)}\cdot\left[\log(4M_{\rho})-\log(1-\rho^{-2})-\log(\epsilon)\right].$ (20)
One may observe that the two bounds only differs from the term
$m^{*}=m_{\text{asym}}-m_{\text{sym}}=\log(1+\rho^{-1})/2\log(\rho).$
Corollary 4.3 studies the lower and upper bounds of such difference.
Corollary 4.3.
Let $A\in\mathbb{R}^{n\times n}$ be a real symmetric positive definite matrix with eigenvalues in $[\lambda_{\min},\lambda_{\max}]$, $f$ be an analytic function on $[\lambda_{\min},\lambda_{\max}]$, $\boldsymbol{u}$ be a normalized vector, $\kappa=\lambda_{\max}/\lambda_{\min}$ be the condition number. Then the difference $m^{*}$ of iterations in the cases of asymmetric and symmetric quadrature nodes should satisfy
$\frac{(\sqrt{\kappa}-1)^{2}}{8\sqrt{\kappa}}\leq m^{*}\leq\frac{\sqrt{\kappa}-1}{4}$ (21)
to guarantee the same computation precision of $Q_{m}(\boldsymbol{u},f,A)$.
Proof 4.4.
As mentioned, the difference between (19) and (20) lies in the term $m^{*}=\log(1+\rho^{-1})/2\log(\rho)$, so the inequalities (21) result from analyzing the lower and upper bounds of it. Note that
$1-\frac{1}{x}\leq\log(x)\leq x-1,\text{ for all }x>0,$
$\frac{x}{1+x}\leq\log(1+x)\leq x,\text{ for all }x>-1.$
Since $\rho=(\sqrt{\kappa}+1)/(\sqrt{\kappa}-1)>1$, the lower bound becomes
$\frac{\log(1+\rho^{-1})}{2\log(\rho)}\geq\frac{\frac{1}{\rho+1}}{2(\rho-1)}=\frac{1}{2(\rho^{2}-1)}=\frac{(\sqrt{\kappa}-1)^{2}}{8\sqrt{\kappa}},$
and the upper bound reads
$\frac{\log(1+\rho^{-1})}{2\log(\rho)}\leq\frac{\frac{1}{\rho}}{2(1-\frac{1}{\rho})}=\frac{1}{2(\rho-1)}=\frac{\sqrt{\kappa}-1}{4},$
which finishes the proof.
One may observe that when matrix becomes ill-conditioned (i.e., $\kappa$ becomes larger), the difference between two bounds of $m$ in asymmetric and symmetric cases increases. We make further discussions on Corollary 4.3 in Section 5.2.
5 Numerical experiments
5.1 Test on Theorem 3.16
In this section we conduct experiments to verify whether Theorem 3.16 is valid. The first three cases study the reproducible matrix $A=H\Lambda H^{T}\in\mathbb{R}^{n\times n}$ with the diagonal matrix
$\Lambda=\mathrm{diag}(\lambda_{1},\lambda_{2},\cdots,\lambda_{n}),$
and the Householder matrix is constructed according to [51]
$H=I_{n}-\frac{2}{n}(\mathbf{1}_{n}\mathbf{1}^{T}_{n}),$
where $I_{n}$ represents the identity matrix of size $n$ and $\mathbf{1}_{n}$ denotes an $n$-dimensional vector of all-ones. The eigenvalues stored in $\Lambda$ are predetermined and $n=50$ is fixed. The fourth case focuses on the nd3k matrix from [15].
    Case 1: $\{\lambda_{i}\}_{i=1}^{50}=\{i/50\}_{i=1}^{50}$, $\boldsymbol{v}=\mathbf{1}_{50}/\sqrt{50}$;
    Case 2: $\{\lambda_{i}\}_{i=1}^{50}=\{1/(51-i)\}_{i=1}^{50}$, $\boldsymbol{v}=\mathbf{1}_{50}/\sqrt{50}$;
    Case 3: $\{\lambda_{i}\}_{i=1}^{50}=\{i/50\}_{i=1}^{50}$, $\boldsymbol{v}=(1,2,\cdots,50)^{T}/\Arrowvert(1,2,\cdots,50)^{T}\Arrowvert$;
    Case 4: nd3k matrix, $\boldsymbol{v}=(1,\cdots,1,-1,\cdots,-1)^{T}/\sqrt{9000}\in\mathbb{R}^{9000}.$
Details of these 4 cases can be seen in Table 1.
Table 1: Details of 4 cases
Case 1 Case 2 Case 3 Case 4
Do $A$ have symmetric eigenvalues? Yes No Yes No
Is $\boldsymbol{\mu}_{1}$ an absolute palindrome? Yes Yes No No
Are Ritz values symmetric? Yes No No No
Any sufficient conditions? Yes ? ? ?
Recall that the figure of $\mu(t)$ would be central symmetric about $(\bar{\lambda},\mu(\bar{\lambda}))$ if $A$ has symmetric eigenvalue distribution and $\boldsymbol{\mu}_{1}$ is an absolute palindrome. Figure 1 showcases the measure function $\mu(t)$ alongside the corresponding $10$ Ritz values for the four cases, offering a visual depiction that substantiates the validity of Theorem 3.16 to a certain degree.
Indeed, it is imperative to acknowledge that while Theorem 3.16 serves as a sufficient condition guaranteeing the symmetry of Ritz values produced by the Lanczos iteration, it is not strictly necessary. Researchers may delve deeper to explore whether alternative conditions exist that render these Ritz values symmetric in cases $2,3,4$.
(a) Case 1: symmetric Ritz values
(b) Case 2: asymmetric Ritz values
(c) Case 3: asymmetric Ritz values
(d) Case 4: asymmetric Ritz values
Figure 1: Plots of discrete measure functions (4) and the locations of Ritz values in four cases.
5.2 Test on Corollary 4.3
We conduct another test on the difference $m^{*}$ between the numbers of Lanczos iterations that ensure the same precision of approximating quadratic forms (1) based on asymmetric and symmetric quadrature nodes. Fig. 2 depicts the lower and upper bounds of $m^{*}$, and the average of two bounds along with the exact $m^{*}$ for different condition numbers $\kappa=10,50,100,500,\ldots,5\times 10^{4},10^{5}$. The elliptical radius is chosen as $\rho=(\sqrt{\kappa}+1)/(\sqrt{\kappa}-1)$.
Figure 2: Lower, upper bounds, the average of the two bounds, and the exact values of the absolute difference $m^{*}$ between the Lanczos iterations required for equivalent accuracy in approximating quadratic forms (1) with asymmetric and symmetric quadrature nodes, where $\kappa=10,50,100,500,\ldots,5\times 10^{4},10^{5}$.
It is observed that the exact difference $m^{*}$ can be roughly substituted by the average of its lower and upper bounds in (21). When $\kappa$ becomes large, e.g., $\kappa=10^{5}$, the gap of Lanczos iterations is around $55$, which leads to huge computational costs for extremely large matrices.
6 Concluding remarks
This paper stems from discrepancies in the analyses of quadrature error observed between [48] and [14]. The difference results from the symmetry of quadrature nodes (Ritz values) in the Lanczos quadrature method. Consequently, we introduce and establish a sufficient condition ensuring this symmetry. Furthermore, for applications involving the estimation of quadratic forms, we investigate the disparity between the minimum Lanczos iterations required to guarantee equivalent approximation accuracy in cases of asymmetric and symmetric quadrature nodes.
We expect the emergence of other theories on the symmetry of Ritz values in the Lanczos method, e.g., sufficient and necessary conditions of such symmetry. We believe that based on such elucidations one may correctly develop further analyses on the Lanczos quadrature method [26], the properties of Ritz values and facilitate the study of other Gauss quadrature rules [49].